What is the molality of a solution containing 15.0 g of NaCl dissolved in 1000.0 g of water?
1
Expert's answer
2018-03-29T07:57:03-0400
Answer:

n(NaOH) = m/M

where n is amount of substance, mole; M is molar mass, g/mole, m is mass of substance, g

n(NaOH) = 10/40 = 0.25 mole

m(H₂O) = 1 kg H₂O

molality = moles of solute/kilograms of solvent

molality = 0.25/1 = 0.25 mole/kg